# Large Angular Momentum

### v1

@@ 3 Quantum and classical angular momentum states
For these state, it makes sense to interpret them “classically” as angular momentum pointed towards ±$\pm{\hat{z}}$, as its fluctuations $(\sim\mathcal{O}(\sqrt{j})$) become negligible with respect to its magnitude $j$, in the limit $j\to\infty$. The same holds true in all states in which spin is oriented towards a generic direction ${\mathbf{n}}$, as has been already noted in the previous section, Eqs. (2.16)-(2.18). Accordingly, we propose the correspondence (by setting $\hbar=1$)
$\boxed{\quad|j,{\mathbf{n}}\rangle\quad\longleftrightarrow\quad{\bf j}^{(\rm class)}=j\,{\bf n}\;\quad}$ (3.11)
@@
${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;.$ (3.12)
Actually, all states “almost oriented towards ${\mathbf{n}}$”, $|j,{\mathbf{n}},q\rangle,$
@@
$\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,,\,\,\,{\rm finite},$ (3.13)
discussed in Sec. 2.2.3 (see (2.22)), can also be interpreted as ${\bf j}^{(\rm class)}=j\,{\bf n}$. The aim of our present investigation is to examine the consistency of such a picture, with respect to the SG processes (this has already been done in Sec. 2.2), the addition rule of the angular momenta (Sec. 4 below) and the rotation matrix (Sec. 5 below).
As for the states with strong quantum fluctuations $|j,m\rangle$, (3.9) (with $m$ generic and not close to ±$\pm j$), they are certainly quantum mechanical, independently of $j$. The same can be said of a generic spin $j$ wave function, (3.4). If such a particle is sent into a SG set-up, it could split in an arbitrary distribution of sub wavepackets, in contrast to the cases discussed in Sec. 2.2. This results in a macroscopic quantum system. It is possible that such a strongly fluctuating, macroscopic quantum angular-momentum state can be realized experimentally, if the whole system is brought to extremely low temperatures near $T=0$  [^6] We recall that any system is in its quantum-mechanical ground state at $T=0$. Accordingly many macroscopic systems can be brought to a quantum-mechanical state, if sufficiently low temperatures can be attained. See the references cited in [1] for the recent efforts to fabricate such systems experimentally..
To see what happens at finite temperatures (either the body-temperature of the macroscopic body, or a warm environment) instead, it is useful to recall the so-called Born-Einstein dispute and its resolution (see [1]). Einstein strongly rebuked the idea by Born, that the absence of quantum diffusion is sufficient to explain the classical nature (the unique trajectoriy) of a macroscopic body, arguing that a doubly or multiply split wave packet, with their centers separated by macroscopic distances, are allowed by Schrödinger equations even for a macroscopic particle. The latter is certainly non-classical. The missing piece for solving this puzzle turns out to be the temperature [1]. Even though at exceptionally low temperatures such a state is possible, it is not so at finite temperatures. Emission of photons and the ensuing self-decoherence (in the case of an isolated body) [1] or an environment-induced decoherence (for an open system) [5, 6, 7] makes a split system a mixture (a mixed state) essentially instantaneously. Also, they cannot be prepared experimentally, e.g., by a passage through a double slit [1]. A (macroscopic) particle passes either through one or the other slit, due to absence of diffusion.
A similar consideration applies to the general angular momentum states with strong fluctuations, (3.9) or (3.4). In the large $j$ limit, such a state necessarily involves large space extension: a macroscopic quantum state. A macroscopic pure quantum state can be realized only at exceedingly low temperatures  [^7] To tie up possible loose ends of the argument, let us note that a single, spin $1/2$ isolated atom in its ground state (a microscopic system) travelling in a good vacuum, has effectively $T=0$. Thus the fact that its split wave packets can attain a macroscopic extension - i.e. its being in a macroscopic quantum state - is perfectly consistent. To realize a macroscopic system (made of many atoms and molecules) in a pure quantum state is another story: it is much more difficult to prepare the necessary low temperatures close to $T=0$. . See [1] for more discussions. At finite temperatures, macroscopically split wave packet of such a particle under the SG set-up is necessarily an incoherent mixture.
What is rather remarkable, perhaps, is the fact that the angular momentum states of minimum fluctuations, (3.12), (3.13), though pure, effectively become classical at $j\to\infty$, without need of any (thermal or environment-induced) decoherence effects, hence independently of the temperatures. This is in agreement with the known general semi-classical behavior of a (pure-state) wave function in QM in the limit $\hbar\to 0$.
3.1 Orbital angular momentum

### v3

@@ 2 Space of quantum spin states and classical angular momenta
For these state, it makes sense to interpret them as a classical angular momentum pointed towards ±$\pm{\hat{z}}$, as its transverse fluctuations $(\sim\mathcal{O}(\sqrt{j})$) are negligible with respect to its magnitude $j$, in the limit $j\to\infty$. The same can be said of the states of spin almost oriented towards $\hat{z}$, $|j,\pm(j-r)\rangle$, $r\ll j$.
Naturally, all states of the form, $|j,{\mathbf{n}}\rangle$, in which spin is oriented towards a generic direction ${\mathbf{n}}$,
${\hat{J}}^{2}|j,{\mathbf{n}}\rangle=j(j+1)|j,{\mathbf{n}}\rangle\;,\qquad\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}}\rangle=j\,|j,{\mathbf{n}}\rangle\;,$ (2.15)
have this property: its fluctuations in the components perpendicular to ${\mathbf{n}}$ become negligible in the $j\to\infty$ limit. The states (2.15) are known as Bloch states or coherent spin states in the literature [5, 10, 6, 11, 7].
These observations lead us to propose that the quantum-classical correspondence (by setting $\hbar=1$) to be made is
$|j,{\mathbf{n}}\rangle\quad\longleftrightarrow\quad{\bf j}^{(\rm class)}=j\,{\bf n}\;.$ (2.16)
Of course, all states “almost oriented towards ${\mathbf{n}}$”, $|j,{\mathbf{n}},q\rangle,$
$\left({\hat{\mathbf{J}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,,\,\quad q\ll j$ (2.17)
discussed in Sec. 3.2.2 (see Eq.(3.17)) below, can also be interpreted as ${\bf j}^{(\rm class)}=j\,{\bf n}$. We are going to examine below the consistency of such a picture, with respect to the SG processes (Sec. 3 below), a study of orbital angular momentum (Sec. 4), the addition rule of the angular momenta (Sec. 5) and the rotation matrix (Sec. 6).
A powerful idea, useful both for mathematical and physical reasoning below, is that the general spin $j$ can be regarded as a direct product state of $n=2j$ spins, each carrying 2Planck-constant-over-$\frac{\hbar}{2}$. From a group-theoretic point of view, it is well known that any irreducible representation of the SU(2) group, which is the double covering group (see e.g., [25]) of the rotation group SO(3), is associated with the Young tableaux,
[…]
with a single row of $n$ boxes, i.e., the totally symmetric direct product of $n=2j$ spin $\frac{1}{2}$ objects. Indeed, the entire theory of quantum angular momentum can be reconstructed based on this point of view [26], by using the operators,
$[a_{i},a^{\dagger}_{j}]=\delta_{ij}\;.$ (2.20)
[…]
Physically, on the other hand, one may consider (2.18) as a particular state of a toy-model macroscopic body made of $n$ atoms, each carrying spin $\frac{1}{2}$.
Now a particularly attractive picture of the quantum states (2.18) of general spin $j$, follows from the so-called stellar representation of $\mathbf{CP}^{n}$ ($n=2j$) [24]. Let the complex numbers
$(Z_{0},Z_{1},\ldots,Z_{n})\sim\lambda\,(Z_{0},Z_{1},\ldots,Z_{n})\;,\qquad\lambda\in{\mathbf{C}}/{\{0\}}$ (2.21)
be the homogeneous coordinates of $\mathbf{CP}^{n}$, and consider the $n$-th order equation,
$\sum_{k=0}^{n}Z_{k}\,z^{n-k}=Z_{0}\,\prod_{\ell=1}^{n}(z-w_{\ell})=0\;.$ (2.22)
As the equation (2.22) is invariant under the rescaling (2.21), the collection of the $n$ roots for $z$, $\{w_{\ell}\}$, can each be regarded as equivalent to a point of $\mathbf{CP}^{n}$. Note that no generality is lost by considering the neighborhood $Z_{0}\neq 0$ above: for instance, one may introduce the variable $z^{\prime}\equiv 1/z$ and rewrite equation (2.22) in terms of $z^{\prime}$. One sees then that the equation with $Z_{0}=0$ contains (at least one) root $z^{\prime}=0$, that is $z=\infty$.
The explicit relation between $\{w_{\ell}\}$ and the $\mathbf{CP}^{n}$ point can be found by working backward: given any collection of $n$ roots $\{w_{\ell}\}$ in the complex plane, one can find the coefficients $\{Z_{k}\}$ the coordinates of $\mathbf{CP}^{n}$ by simply expanding the second expression of (2.22). The connection between this construction and the picture (2.18) above comes from the fact that any complex number $\{w_{\ell}\}$ can be regarded as the stereographic projection of a point on a sphere $S^{2}$, from the north (or south) pole, onto the plane containing the equator. See Fig. 2. Naturally, a point on $S^{2}$ can be identified with the unit vector ${\mathbf{n}}$, (2.3), the orientation of each component spin $\frac{1}{2}$. The explicit relation is
$w_{\mathbf{n}}=\frac{e^{i\phi}}{\tan(\theta/2)}\;.$ (2.23)
Collecting things, we see that a point in $\mathbf{CP}^{n}$ is represented by an unordered collection of $n$ points on a sphere (stars), see Fig. 3. This is known as the stellar representation of $\mathbf{CP}^{n}$.
Figure 2: Two points (stars) on the sphere, and their stereographic projections from the north pole (N.P.) onto the plane containing the equator.
Figure 3: A general point in $\mathbf{CP}^{n}$ is equivalent to an unordered collection of $n$ points (stars) on a sphere, in the stellar representation, (2.22). Each star is associated with the complex number $w_{\ell}$ by a stereographic projection, as in Fig. 2.
From the group-theoretic point of view, (2.18), a generic quantum state of spin $j$, is equivalent to the collection (the direct product) of $n=2j$ spin $\frac{1}{2}s,$
${\cal S}\,\,|{\bf n}_{1}\rangle|{\bf n}_{2}\rangle\ldots|{\bf n}_{n}\rangle$ (2.24)
where ${\cal S}$ stands for a symmetrization, orientated towards various directions ${\mathbf{n}}_{k}$, $k=1,2,\ldots,n$, see Eq.(2.5). (2.24) represents a general state of spin $j=2n$ and naturally corresponds to the random sets of stars in Fig. 3.
The particular state in which the spin $j$ is oriented towards a definite direction ${\mathbf{n}}$, (2.15), corresponds to the case in which the component $\frac{1}{2}$ spins are all oriented towards the same direction ${\mathbf{n}}_{1}={\mathbf{n}}_{2}=\ldots={\mathbf{n}}_{n}={\mathbf{n}}$,
$|j,{\mathbf{n}}\rangle=|{\bf n}\rangle|{\bf n}\rangle\ldots|{\bf n}\rangle$ (2.25)
(which is automatically symmetric). See Fig. 4. By using (2.5) for each $|{\bf n}\rangle$, and collecting terms with the same fixed number $k$ of the spin up factors, we find $|j,{\mathbf{n}}\rangle$ as an expansion in terms of the eigenstates of $J_{z}$ with $J_{z}=m$,
$|j,{\mathbf{n}}\rangle=\sum_{k=0}^{2j}c_{k}\,|j,m\rangle\;,\qquad m=-j+k\;,$ (2.26)
$c_{k}={\binom{2j}{k}}^{1/2}\,e^{i(j-k)\phi}\left(\cos\tfrac{\theta}{2}\right)^{k}\left(\sin\tfrac{\theta}{2}\right)^{2j-k}\;,\qquad\sum_{k=0}^{2j}|c_{k}|^{2}=1\;,$ (2.27)
where $\binom{n}{k}=n!/k!(n-k)!$ are the binomial coefficients. The states (2.26), (2.27), coincide, apart from an overall phase factor, with the Bloch (or coherent spin) states defined in [5, 10, 6, 11, 7], and denoted as $|j,\Omega\rangle$, $|j,\{\theta,\phi\}\rangle$, etc.
In order to understand this formula from the point of view of the stellar representation of $\mathbf{CP}^{n}$, we need to translate the homogeneous coordinates of $\mathbf{CP}^{n}$ to the coefficients $\{c_{k}\}$ in (2.26) as
$Z_{k}=c_{k}^{2}\;,\qquad k=0,1,\ldots,n\;,$ (2.28)
and at the same time replace the projection of the stars on $S^{2}$ as
$w_{\ell}\to w_{\ell}^{2}\;,\qquad\forall\ell\;.$ (2.29)
Eq.(2.28) and (2.29) reflect the double covering of SO(3) group by SU(2).
In the stellar representation of $\mathbf{CP}^{n}$, (2.21), (2.22), the state (2.25) corresponds to the situation where all roots coincide,
$w_{\ell}=w_{\mathbf{n}}\;,\qquad\forall\ell\,.$ (2.30)
All stars are at a coincident point ${\mathbf{n}}$ on the sphere, see Fig. 4.
Figure 4: The coherent spin state (2.15), (2.26), (2.27) in the stellar representation.
Taking into account the doubling (2.28) and (2.29) and expanding ($n=2j$)
$\sum_{k=0}^{n}(c_{k})^{2}\,z^{n-k}=\prod_{\ell=1}^{n}(z-w_{\ell}^{2})=(z-w_{\mathbf{n}}^{2})^{n}\;,$ (2.31)
($w_{\mathbf{n}}=\frac{e^{i\phi}}{\tan(\theta/2)}$), one finds precisely (2.26), (2.27), apart from unobservable phases and an overall factor.
Thus the stellar representation of $\mathbf{CP}^{n}$ points - the generic states of spin $j$ provides us with an appealing intuitive picture of quantum fluctuations. The $\mathbf{CP}^{n}$ points with randomly distributed stars, Fig. 3, represent spin states with large quantum fluctuations. The special states, (2.15), (2.25) - (2.27) with coincident stars, Fig. 4, instead, are those with the minimum fluctuations, and correspond to the Bloch or spin coherent state [5, 10, 6, 11, 7]. In the next four sections we shall illustrate how these latter states effectively behave as the classical angular momenta, $j\,{\mathbf{n}}$, in the $j\to\infty$ limit.
3 Stern-Gerlach experiments
In this section we compare the Stern-Gerlach experiment [27] for small and large spins, with special attention to the Bloch states, (2.15), (2.26), (2.27).
3.1 Spin $1/2$ atom
For definiteness let us take an incident atom with spin $j=\tfrac{1}{2}$ directed along a definite, but generic, direction ${\bf n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, i.e.
$\Psi=\psi({\bf r},t)\,|{\bf n}\rangle\;,$ (3.1)
Let us suppose that the atom enters from $x=-\infty$ and moves towards $x=+\infty$. When the atom passes through a region with an inhomogeneous magnetic field with
$\frac{\partial B_{z}}{\partial z}\neq 0\;,$ (3.3)
the wave packet splits into two subpackets, with respective weights, $|c_{1}|^{2}$ and $|c_{2}|^{2}$: they manifest as the relative intensities of the two separate image bands on the screen at the end.
[…]
The original Stern-Gerlach experiment (1922) [27] has had a fundamental impact to the development of quantum theory, in proving the quantization of the angular momentum, the existence of half-integer spin, and more generally, demonstrating the quantum-mechanical nature of the silver atom as a whole [23].
The wave function of the form (3.2) corresponds to a pure state, i.e., $100$% polarized beam where all incident atoms are in the same spin state. If the beam is partially polarized or unpolarized, the state is described by a generic density matrix $\rho$, with
$\mathop{Tr}\nolimits\,\rho=1\;,\qquad\rho_{ii}\geq 0\;,\qquad i=1,2\;.$ (3.4)
The Stern-Gerlach experiment measures the relative frequency that the atom arriving at the screen happens to have spin $s_{z}=\tfrac{1}{2}$ or spin $s_{z}=-\tfrac{1}{2}$. Let
$\Pi_{\uparrow}=|\uparrow\rangle\langle\uparrow|=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right)\;,\qquad\Pi_{\downarrow}=|\downarrow\rangle\langle\downarrow|=\left(\begin{array}[]{cc}0&0\\ 0&1\end{array}\right)$ (3.5)
be the projection operators on the spin up (down) states; the relative intensities of the upper and lower blots on the screen are, according to QM,
${\bar{\Pi}_{\uparrow}}=\mathop{Tr}\nolimits\Pi_{\uparrow}\rho=\rho_{11}\;;\quad{\bar{\Pi}_{\downarrow}}=\mathop{Tr}\nolimits\Pi_{\downarrow}\rho=\rho_{22}\;,\qquad\rho_{11}\ +\rho_{22}=1\;.$ (3.6)
The prediction about the relative intensities of the two narrow atomic image bands from the wave function (3.2) and from the density matrix, (3.6), is in general indistinguishable, as is well known.
3.2 Large spin
Consider now the state of a spin $j$ directed towards a direction ${\mathbf{n}}$, $|j,{\mathbf{n}}\rangle$, given in (2.15),(2.26), and (2.27), Fig. 4. Choosing the magnetic field (and its gradients) in the ${z}$-direction, we need to express $|j,{\mathbf{n}}\rangle$ as a superposition of the eigenstates of ${\hat{J_{z}}}$ ($j=n/2$). This is precisely the expression given in Eq.(2.26) and Eq.(2.27).
Using Stirling’s formula in (2.27), one finds, for $n$ and $k=j+m$ both large with $x=k/n$ fixed, the following distribution in different values of $m=-j+k$,
$|c_{k}|^{2}\simeq e^{nf(x)}\;,\qquad x=k/n=(j+m)/2j\;,$ (3.7)
@@ 3.2 Large spin
The saddle-point approximation valid at $n\to\infty$, yields
and therefore
[…]
$J_{z}=m=n(x-\tfrac{1}{2})=j\,(2\cos^{2}\tfrac{\theta}{2}-1)=j\,\cos\theta\;.$ (3.11)
This means that a large spin ($j\gg\hbar$) quantum particle with spin directed towards ${\mathbf{n}}$, in a Stern-Gerlach setting with an inhomogeneous magnetic field (3.3), moves along a single trajectory of a classical particle with $J_{z}=j\,\cos\theta$, instead of spreading over a wide range of split sub-packet trajectories covering $-j\leq m\leq j$. See Fig. 5 (spin $j=10^{3}$) and Fig. 6 ($j=2\cdot 10^{5}$) and compare them with Fig. 1 for spin $j=13/2$.
Figure 5: The distribution $|c_{k}|^{2}$ in $k$ for possible values of $J_{z}=m$ ($m=-j+k$), for a spin $j=1\cdot 10^{3}$ particle in the state (2.26), (2.27), with $\theta=\pi/2$ (center, orange peak) or with $\theta=\pi/4$ (right, blue peak). The particle starts looking like classical.
Figure 6: The distribution $|c_{k}|^{2}$ ($J_{z}=m=-j+k$) as in Fig. 5 but for a spin $j=2\cdot 10^{5}$, for $\theta=\pi/2$ (center, orange peak) or $\theta=\pi/4$ (right, blue peak). This figure is drawn by using the approximation (3.7) and (3.8), rather than the exact formula, (2.26), (2.27).
This (perhaps) somewhat surprising result appears to indicate that quantum mechanics (QM) takes care of itself, so to speak, in ensuring that a large spin particle ($j/\hbar\to\infty$) behaves classically, at least for these particular states $|j,{\mathbf{n}}\rangle$. No extra conditions are necessary. This is consistent with the known general behavior of the wave function in the semi-classical limit ($\hbar\to 0$)  [^5] Of course this does not mean that the classical limit necessarily requires or implies $j/\hbar\to\infty$.. Note that if the value $j\gg\hbar$ is understood as due to the large number of spin $\frac{1}{2}$ particles composing it (see the comment after (2.27)), the spike (3.9), (3.10) can be understood as due to the accumulation of a huge number of microstates giving $J_{z}\simeq m$.
Another, equivalent way to see the shrinkage of the distribution in $J_{z}=m$, (3.10), is to study its dispersion. Given the relative weights $|c_{k}|^{2}$, (2.27), where $m=-j+k$ and $j=\tfrac{n}{2}$, with $k=0,1,\ldots,n$, it is a simple exercise to find that
$\langle J_{z}\rangle=j\,\cos\theta\;,\qquad\langle J_{z}^{2}\rangle=j^{2}\cos^{2}\theta+\frac{j}{2}\sin^{2}\theta\;,$ (3.12)
so that the dispersion (the fluctuation width) is given by  [^6] These results are well known [10].
$(\Delta J_{z})^{2}=\langle(J_{z}-\langle J_{z}\rangle)^{2}\rangle=\langle J_{z}^{2}\rangle-\langle J_{z}\rangle^{2}=\frac{j}{2}\sin^{2}\theta\;.$ (3.13)
It follows that
$\lim_{j/\hbar\to\infty}\frac{|\Delta J_{z}|}{j}=0\;,$ (3.14)
i.e. it is an infinitely narrow distribution. Note that the SG experiment is indeed a measurement of $J_{z}$ in the state (2.26). The fact that the result is always $J_{z}\sim\langle J_{z}\rangle=j\,\cos\theta$ with no dispersion, means that it is a classical angular momentum $\sim j\,{\mathbf{n}}$.
3.2.1 The state preparation and successive SG set-ups
For a small $j$, as it is well known, the SG set-up can be used to prepare any desired state $|j,J_{z}=m\rangle$, by an intentionally biased SG apparatus, e.g. in which all the split sub-wavepackets are blocked, except for the one corresponding to $|j,m\rangle$. Equivalently, one may use the so-called null-measurements (see Ref. [30] for a recent review). In any case, for a small finite $j$, where there are $12j+$ 1 well-separated sub-packets (see Fig. 1), there is no difficulty in extracting any desired quantum state for subsequent experiments (state preparation).
In the large-spin limit ($j\to\infty$) of the Bloch states $|j,{\mathbf{n}}\rangle$ we are considering, there is no possibility of extracting or selecting a desired sub-wavepacket, since there is essentially only one narrow wavepacket present (see (3.11), Fig. 6). Trying to extract a state with $J_{z}$ not close to $j\cos\theta$ would simply give a zero result. We have in mind a macroscopic scenario where, in comparison to the finite region (in principle) occupied by the waves $-j\leq m\leq j$, the fluctuations of the values of $m$ around (3.11) are of the order of $\sqrt{j}\ll j$, i.e. of zero width.
We might consider introducing a second set of Stern-Gerlach (SG) magnetic fields directed in a different direction, denoted by ${\hat{z}}^{\prime}$, as a new quantization axis. This setup allows us to study the state emerging from the first set of SG magnetic fields (but without an imaging screen),
$|j,{\mathbf{n}}\rangle\sim c_{k}\,|j,J_{z}=m\rangle\;,\qquad m\sim j\,\cos\theta\;.$ (3.15)
Since the sum of states around $m\sim j\,\cos\theta$ is essentially equal to the complete sum, this state can be rewritten as
$|j,{\mathbf{n}}\rangle$ $=$ $\sum_{m,m^{\prime}}c_{k}\,|j,m^{\prime}\rangle\langle j,m^{\prime}|j,m\rangle$ (3.16)
$=$ $\sum_{m,m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|j,m\rangle\langle j,m|j,{\mathbf{n}}\rangle$
$=$ $\sum_{m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|j,{\mathbf{n}}\rangle=\sum_{m^{\prime}}c_{k^{\prime}}|j,m^{\prime}\rangle\;,$
where in the first line a completeness relation $\sum_{m^{\prime}}|j,m^{\prime}\rangle\langle j,m^{\prime}|={\mathbf{1}}$ has been inserted, and in the second we recalled $c_{k}=\langle j,m|j,{\mathbf{n}}\rangle$ (see Eq.(2.26)). (3.16) represents an expansion of the original state $|j,{\mathbf{n}}\rangle$ as a linear combination of the eigenstates of $J_{z^{\prime}}=m^{\prime}$. In the limit $j\to\infty$, the results from Section 3.2 tell us that the sum is dominated by the small region around $m^{\prime}\sim j\cos({\mathbf{n}}\cdot{\hat{z}}^{\prime})$, with negligible fluctuations. The trajectory remains unique, regardless of the choice of quantization axis (i.e. the direction of the magnetic field) ${\hat{z}}^{\prime}$.
3.2.2 States almost directed towards $\vec{\mathbf{n}}$
Obviously, in the large spin limit, other states close to $|j,{\mathbf{n}}\rangle$,
@@ 3.2.2 States almost directed towards $\vec{\mathbf{n}}$
$\left({\hat{\mathbf{j}}}\cdot{\mathbf{n}}\right)\,|j,{\mathbf{n}},q\rangle=(j-q)\,|j,{\mathbf{n}},q\rangle\;,\qquad q=1,2,\ldots\,\,\,{\rm fixed}$ (3.17)
$|j,{\mathbf{n}},1\rangle$ $=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}\,|{\bf n}\rangle_{1}\otimes\cdots|{\bf n}\rangle_{\ell}\otimes|-{\bf n}\rangle\otimes|{\bf n}\rangle_{\ell+2}\otimes\cdots|{\bf n}\rangle_{N}\;,$ (3.18)
$=$ $\sum_{\ell=0}^{N-1}\frac{1}{\sqrt{N}}|j_{1},{\mathbf{n}}\rangle\otimes|-{\bf n}\rangle\otimes|j_{2},{\mathbf{n}}\rangle\;,$
where $|j_{1},{\mathbf{n}}\rangle$ and $|j_{2},{\mathbf{n}}\rangle$ ($j_{1}=\tfrac{\ell}{2}\;,j_{2}=\tfrac{N-\ell-1}{2}$), are two spins oriented towards ${\mathbf{n}}$, (2.15), as already studied in Sec. 3.2. In the $j\to\infty$ limit, the sum is dominated by terms with $\ell\sim\mathcal{O}(N)\to\infty$. The SG projection of this state onto the $j_{z}$ eigenstates is therefore approximated by
$j_{z}\sim j_{1\,z}+j_{2\,z}=j_{1}\cos\theta+j_{2}\cos\theta\simeq j\,\cos\theta\;,$ (3.19)
where the results of Sec. 3.2 have been used. A similar result holds for $|j,{\mathbf{n}},q\rangle$ for any finite $q$, showing that all these states behave as a classical angular momentum, $j\,{\mathbf{n}}$.
4 Orbital angular momentum